OpenStudy (anonymous):

if swimming pool is 1500 meters and the length is 10 meter minus twice the width.. what is the width?

OpenStudy (bahrom7893):

A = 1500m = LW L = 10 - 2W

OpenStudy (bahrom7893):

A = 1500m = (10-2W)*W

OpenStudy (bahrom7893):

1500 = 10W - 2W^2 2W^2-10W+1500 = 0

OpenStudy (anonymous):

have that but what would actually width be then?

OpenStudy (bahrom7893):

weird, is the area 1500 meters^2 or the perimeter 1500 meters?

OpenStudy (bahrom7893):

cuz u can't solve for W

OpenStudy (bahrom7893):

if area is given...